OpenStudy (anonymous):

if the area of a rectangle is 24a^2b feet in the length is 8ab^2 what would be the width of the rectangle given that is found by dividing area by length simplify the answer? A. 3a/b B. B/3a C. 3/ab D. 3ab NEED HELP!!!!!!!!😓😓😓😓😓

OpenStudy (taramgrant0543664):

So we can break this down kind of into smaller sections to make it easier So 24/8=3 a^2/a is the same as a^(2-1)=a b/b^2 is the same as b^(1-2)=b^(-1) or 1/b So we can then put it all together as: 3a/b

OpenStudy (anonymous):

im sorry i cant help

OpenStudy (anonymous):

Thank you!! That helped me a lot!! Just breaking it down helps!

OpenStudy (taramgrant0543664):

No problem! Breaking things down into small steps helps a lot sometimes!!